# ShowHowTo: Generating Scene-Conditioned Step-by-Step Visual Instructions

### v1

@@ 1 Introduction
This paper makes the following contributions. (1) We introduce the problem of generating a sequence of visual instructions conditioned on an input image. (2) We introduce a fully automatic approach for collecting step-by-step visual instruction training data from in-the-wild instructional videos, creating a large-scale, high-quality dataset of 0.6M step-by-step instruction sequences of 4.5M image-text pairs. (3) We train a video diffusion model capable of generating sparse step-by-step visual instructions consistent with the provided input image. (4) We evaluate our generated sequences across three aspects (step, scene, and task) and show our model achieves state-of-the-art results on all of them.
@@ 3.1 Automatic Dataset Collection
Our approach takes as input a narrated instructional video for a specific task. First, it extracts a sequence of key steps in the form of concise, free-from textual instructions from the video’s narration. Then, it associates each step with the corresponding keyframe in the video. The output is an ordered sequence of image-text pairs. This is a very challenging task due to the high level of noise, the possible misalignment of the narration and the visual content as well as the sheer variety of visual appearance and the spoken natural language in the input internet videos.
@@
Cross-modal frame alignment. For each instructional step, our goal is to identify a single representative frame that best demonstrates the instruction visually. While contrastive models [47, 66, 21] can align text instructions with frames, we observe that naive text matching to all of the thousands of video frames leads to noisy results. Therefore, we limit the alignment to the identified instruction step temporal interval, expanded by $\epsilon=15$ seconds to allow for some level of misalignment between the narrations and visual demonstrations [25]. Given these expanded intervals, we compute text-frame similarity scores using DFN-CLIP [21] and select the best alignment that satisfies the temporal ordering of the steps. We provide more details about the matching process in Appendix A.
@@ 5 Experiments
We first introduce our evaluation setup in Section 5.1. Then, in Section 5.2, we compare our method with the current state-of-the-art quantitatively on the test set and through a user study. Section 5.3 analyzes key designs of our method through ablation studies. Finally, in Section 5.4, we showcase the qualitative results of our method. For implementation details and additional results, please see the appendix.
@@ 5.1 Evaluation details
Dataset. We construct train and test splits from our dataset of 578K samples. Our test set comprises 3,964 sequences from 200 tasks covering the distribution over task categories in the full dataset. To ensure sample quality, we prioritize samples with high DFN-CLIP alignment scores as measured in our dataset creation pipeline (Section 3).
Evaluation metrics. We evaluate our model by measuring the correctness and consistency of the generated visual instructions using similar metrics as Menon et al. [41]. We describe the used metrics next with full details available in Appendix D. (1) Step Faithfulness [41] measures whether each generated image $\hat{I}_{i}$ correctly depicts its corresponding text instruction $\tau_{i}$. It is computed as the zero-shot accuracy of the DFN-CLIP model where the generated image $\hat{I}_{i}$ is classified into classes $\{\tau_{i}\}_{i=0}^{n}$ of all text instructions of the sequence. (2) Scene Consistency measures whether the generated image $\hat{I}_{i}$ consistently captures the scene from the input image $I_{0}$ (e.g., the same utensils are used on the same kitchen countertop as in the input image). Intuitively, a generated image $\hat{I}_{i}$ is considered scene-consistent if it visually matches any frame from its source video $\{I_{i}\}_{i=1}^{n}$ (excluding the input image to avoid trivial copy solution). Therefore, for each generated image $\hat{I}_{i}$, the most similar image according to the DINOv2 [44] score is retrieved from the test set images. The metric then measures if the retrieved image is from the same video as the input. (3) Task Faithfulness measures how well the generated sequence $\{\hat{I}_{i}\}_{i=1}^{n}$ represents its intended task. It is measured as the zero-shot accuracy of the DFN-CLIP model where the generated sequence’s averaged feature vector is classified into all 200 test set tasks. In contrast to Menon et al. [41], the generated sequences are evaluated holistically rather than per-step, as often steps are not unique to a task (e.g., “knead dough” step appears in both “Make sourdough bread” and “Make pizza” tasks), and the classification is done into all test set tasks rather than a small random subset, providing a more robust evaluation. Lastly, we report Overall score as the average of the three metrics.
@@ 5.2 Comparison with the State-of-the-Art
Method Step Scene Task Overall
Faithf. Consist. Faithf.
(a) InstructPix2Pix [12] 0.25 0.17 0.25 0.22
(b) AURORA [35] 0.25 0.33 0.24 0.27
(c) GenHowTo [53] 0.49 0.13 0.27 0.29
(d) Phung et al. [45] 0.36 0.03 0.38 0.26
(e) StackedDiffusion [41] 0.43 0.02 0.42 0.29
(f) ShowHowTo 0.52 0.34 0.42 0.43
(g) Random 0.19 0.00 0.01 0.07
(h) Stable Diffusion [48$\dagger$ 0.70 0.03 0.44 0.39
(i) Copy of the input image 0.19 0.62 0.39 0.40
(j) Source sequences 0.50 1.00 0.56 0.69
@@
Table 1: Comparison with state-of-the-art on the ShowHowTo test set. Out of all the visual instruction generation methods, our method best follows the input prompts. Also, our method’s outputs are the most consistent with the input image.
Compared methods. We compare ShowHowTo to state-of-the-art methods for visual instruction generation as well as various baselines. For image-to-image methods (a-c), we generate the visual instructions sequence by iteratively using the last generated image as the input for the next step generation. (a) InstructPix2Pix [12] is trained to manipulate input images according to a text prompt by training on synthetic paired image data. In contrast, (b) AURORA [35] is trained on a manually curated dataset of image pairs from videos, while (c) GenHowTo [53] extracts the image pairs for training from instructional videos automatically.
Methods that generate image sequences (d-e) do not accept an input image, therefore, we apply the common input masking approach, where the first denoised frame of the sequence is replaced by the noised ground truth frame in each step of the generation. The (d) Phung et al. [45] method generates consistent sequences of visual instructions by attending to all frames in the sequence in the spatial attention layers. As it is a training-free method, we reimplement it and use it with the Stable Diffusion backbone [48]. (e) StackedDiffusion [41] is trained on WikiHow illustrated image sequences. It generates the image sequence as a single tiled image.
@@
Quantitative results. We show the results of different methods on our test set in Table 1. The methods trained to perform localized edits (a-b) generate outputs fairly consistent with the input image (see the Scene Consistency metric), yet they fail to properly capture the instructions described by the text prompts (evidenced by the Step Faithfulness metric). On the other hand, methods for generating sequences of visual instructions (d-e) model the instructions well according to the input text prompts, but they perform poorly in scene consistency. In contrast, our method (f) generates visual instructions that are consistent with the input scene and correctly capture the action specified by the prompt. Our method generates images that are as faithful to the input textual instructions as the dataset sequences (j) (see the Step Faithfulness metric).
Figure 4: User study results. Win rates of the ShowHowTo method against baselines from pairwise forced decision user evaluations, divided into step, scene, and task. Values larger than 50% indicate ShowHowTo is preferred over the method shown on the right.
User study. We present a user study with 9 participants evaluating sequences from 100 randomly sampled tasks from our test set. Each participant compared 50 ShowHowTo sequences against baselines using three criteria: (1) Step Faithfulness (Which sequence better follows the steps?), (2) Scene Consistency (Which sequence is more likely to come from the same video?), and (3) Task Faithfulness (Which sequence accurately depicts the instructions for the task of [task]?). As shown in Figure 4, ShowHowTo outperforms all baselines. Notably, users preferred our generations over sequences from source videos in 42% of cases for both step and scene metrics, which may be attributed to the instructional videos not showing good views of the step at times and, importantly, the high quality of our generation. Lower task faithfulness scores against the source sequences suggest room for improvement in future methods. More details are given in Appendix D.
@@ 5.3 Ablations
We evaluate the key design decisions of our proposed method, i.e., the model conditioning, the training data, and the variable sequence length during training.
Text conditioning type Step Scene Task Overall
@@
Text model conditioning. We evaluate how different types of text conditioning affect the model performance. For video models, it is common to provide a single text prompt as a form of conditioning. However, visual instructions are substantially different from each other, possibly requiring different conditioning. We construct a single prompt for each sequence by concatenating all step prompts and by summarizing the step prompts using an LLM [19]. In Table 2, we show that our choice of separate prompt per step significantly outperforms both of the single prompt variants. Additionally, we demonstrate that using the free-form step description for $\tau_{0}$ outperforms the fixed prompt ‘an image’.
Model training data Step Scene Task Overall
@@
Training data. In Table 3, we analyze the impact of different training datasets by comparing our dataset with two related instructional datasets of similar scale and task coverage: (i) HowToStep [37], which contains automatically extracted video-text sequences from cooking videos, and (ii) WikiHow-VGSI [62], which consists of manually created image-text sequences from WikiHow articles, where the images primarily consist of digitally drawn illustrations. To train on the HowToStep dataset, we select the middle frame of each video segment as the visual instruction frame. We observe significantly worse performance caused both by the lack of precise instruction frame information as well as very noisy video segments (e.g., the dataset contains ‘Thank you for watching!’ segments which are not instructional). Training on WikiHow-VGSI yields higher Step Faithfulness scores, likely due to the quality of manually matched images and prompts. However, the overall performance remains significantly below our approach, as the model primarily learns from illustrated images resulting in less consistent scene generation. Lastly, we compare the performance of the ShowHowTo model trained only on cooking task videos—a decision that other works make (e.g., Li et al. [37])—with our full training dataset showing our model benefits from scaling the data to the open task domain.
Training sequence length Step Scene Task Overall
@@
Variable sequence length. Instructions are often of variable length, therefore, one of the key model requirements is to support generating image sequences of different lengths. While attention-based architectures allow for any sequence length, the question is how to train such a model. We test training the model on variable sequences of up to 4 frames, 8 frames, and 16 frames. We also train the model on sequences of length 8 only. In Table 4, we show that training on shorter sequences up to 4 frames results in high Scene Consistency but low Step Correctness, while training on sequences up to 16 frames is the opposite. This can be attributed to the fact that short sequences often keep the same background across the whole sequence, encouraging the model to preserve the background at the expense of the prompt. Longer sequences, on the other hand, have more variation of the background, e.g., as the task moves from the counter to the hob. The model is thus less likely to enforce the background during inference. Lastly, we also show that it is important to train always on consecutive sequences of visual instructions. If a subset of visual instructions from a video is sampled randomly (with the temporal ordering preserved), the scene consistency is decreased (Table 4, third row).
@@ 5.4 Qualitative Results
Qualitative results in Figure 1, Figure 6, and Figures 13, 14, and 15 in the appendix demonstrate the key strengths of the ShowHowTo model. The model consistently preserves the scene as well as various objects, tools, and ingredients from the user-provided input image (e.g., pot and vegetables in Figure 6, first row). It dynamically adjusts the viewpoint to emphasize key actions. The model can also introduce plausible task-relevant objects (e.g., knives or bowls) if these objects are not present in the user-provided input image. Notably, the model effectively adapts human poses to demonstrate various object manipulations (e.g., flower arranging in Figure 6, third row).
We also compare ShowHowTo with the related methods qualitatively in Figure 5. The related methods for generating instructional sequences (StackedDiffusion [41] and Phung et al. [45]) fail to preserve the scene from the input image. For example, they generate blue, black, or wooden kitchen countertop instead of the glossy white one from the input image. On the other hand, image-to-image approaches (GenHowTo [53], AURORA [35], and InstructPix2Pix [12]) perform minimal edits and propagate errors through the output image sequence due to the iterative generation (e.g., the persistent floating dough generated by the GenHowTo method).
Limitations. While our method can generate complex step-by-step visual instructions conditioned on the input image, it inherits the limitations of the models it is based on as well as introduces new limitations stemming from the novel source of the training data. The ShowHowTo model can struggle to correctly maintain object states across many frames; for example, it can generate an image with raw meat even if the meat has been cooked already in previous images. Also, while the model often correctly generates common objects from instructional videos, for rare objects, such as electrical components, the model may generate objects in physically impossible configurations. Please, see Figure 16 in the appendix for examples of the failure cases.
@@ 6 Conclusion
This work explores, for the first time, generating environment-specific visual instructions to accomplish a user-defined task. We introduce a fully automated and scalable pipeline to create a dataset of 578K instructional image-text sequences from online videos, without requiring any manual supervision. Using this data, we train ShowHowTo, a model for generating contextualized step-by-step visual instructions. Our experiments demonstrate the ShowHowTo model’s superior ability to generate accurate, scene-consistent instructional steps across various HowTo tasks, outperforming existing methods. We believe this work opens new avenues for personalized guidance in assistive technologies and step-by-step goal generation for robot planning.
Acknowledgements. We acknowledge VSB – Technical University of Ostrava, IT4Innovations National Supercomputing Center, Czech Republic, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (grant ID: 90254). Research at the University of Bristol is supported by EPSRC UMPIRE (EP/T004991/1) and EPSRC PG Visual AI (EP/T028572/1). Prajwal Gatti is partially funded by an uncharitable donation from Adobe Research to the University of Bristol. This research was co-funded by the European Union (ERC FRONTIER, No. 101097822 and ELIAS No. 101120237) and received the support of the EXA4MIND project, funded by the European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreement N° 101092944. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 1 Introduction
This paper makes the following contributions. (1) We introduce the problem of generating a sequence of visual instructions conditioned on an input image. (2) We introduce a fully automatic approach to collect step-by-step visual instruction training data from in-the-wild instructional videos, creating a large-scale, high-quality dataset of 0.6M step-by-step instruction sequences of 4.5M image-text pairs. (3) We train a video diffusion model capable of generating sparse step-by-step visual instructions consistent with the input image. (4) We evaluate our generated sequences across three aspects (step, scene, and task) and show our model achieves state-of-the-art results on all of them.
@@ 3.1 Automatic Dataset Collection
Our approach takes as input a narrated instructional video for a specific task. First, it extracts a sequence of key steps in the form of concise, free-form textual instructions from the video’s narration. Then, it associates each step with the corresponding keyframe in the video. The output is an ordered sequence of image-text pairs. This is a very challenging task due to the high level of noise, the possible misalignment of the narration and the visual content as well as the sheer variety of visual appearance and the spoken natural language in the input internet videos.
@@
Cross-modal frame alignment. For each instructional step, our goal is to identify a single representative frame that best demonstrates the instruction visually. While contrastive models [47, 66, 21] can align text instructions with frames, we observe that naive text matching across thousands of video frames leads to noisy results. Therefore, we limit the alignment to the identified instruction step temporal interval, expanded by $\epsilon=15$ seconds to allow for some level of misalignment between the narrations and visual demonstrations [25]. Given these expanded intervals, we compute text-frame similarity scores using DFN-CLIP [21] and select the best alignment that satisfies the temporal ordering of the steps. We provide more details about the matching process in Appendix A.
@@ 5 Experiments
We first introduce our evaluation setup in Section 5.1. In Section 5.2, we compare our method to current state-of-the-art quantitatively on the test set and through a user study. Section 5.3 analyzes key designs of our method through ablation studies. Finally, Section 5.4, showcases qualitative results of our method. For implementation details and additional results, see the appendix.
@@ 5.1 Evaluation details
Dataset. We construct train and test splits from our dataset of 578K samples. Our test set comprises 3,964 sequences from 200 tasks covering the distribution over task categories in the full dataset. To ensure sample quality, we prioritize samples with high DFN-CLIP alignment scores as measured in our dataset creation pipeline (Section 3). For zero-shot evaluation of our method, we also use a random subset of 1442 non-illustrated instructional sequences [^1] See the project website for the list of selected sequences. from the WikiHow-VGSI dataset [62] as an additional test set.
Evaluation metrics. We evaluate our model by measuring the correctness and consistency of the generated visual instructions using similar metrics as Menon et al. [41]. We describe the used metrics next with full details available in Appendix D. (1) Step Faithfulness [41] measures whether each generated image $\hat{I}_{i}$ correctly depicts its corresponding text instruction $\tau_{i}$. It is computed as the zero-shot accuracy of the DFN-CLIP model where the generated image $\hat{I}_{i}$ is classified into classes $\{\tau_{i}\}_{i=0}^{n}$ of all text instructions of the sequence. (2) Scene Consistency measures whether the generated image $\hat{I}_{i}$ consistently captures the scene from the input image $I_{0}$ (e.g., the same utensils are used on the same kitchen countertop as in the input image). Intuitively, a generated image $\hat{I}_{i}$ is considered scene-consistent if it visually matches any frame from its source video $\{I_{i}\}_{i=1}^{n}$ (excluding the input image to avoid trivial copy solution). Therefore, for each generated image $\hat{I}_{i}$, the most similar image according to the DINOv2 [44] score is retrieved from the test set images. The metric then measures if the retrieved image is from the same video as the input. (3) Task Faithfulness measures how well the generated sequence $\{\hat{I}_{i}\}_{i=1}^{n}$ represents its intended task. It is measured as the zero-shot accuracy of the DFN-CLIP model where the generated sequence’s averaged feature vector is classified into all 200 test set tasks. In contrast to Menon et al. [41], the generated sequences are evaluated holistically rather than per-step, as often steps are not unique to a task (e.g., “knead dough” step appears in both “Make sourdough bread” and “Make pizza” tasks), and the classification is done into all test set tasks rather than a small random subset, providing a more robust evaluation.
@@ 5.2 Comparison with the State-of-the-Art
ShowHowTo WikiHow [62]
Method Step Scene Task Step Scene
Faithf. Consist. Faithf. Faithf. Consist.
(a) InstructPix2Pix [12] 0.25 0.17 0.25 0.32 0.12
(b) AURORA [35] 0.25 0.33 0.24 0.33 0.15
(c) GenHowTo [53] 0.49 0.13 0.27 0.60 0.06
(d) Phung et al. [45] 0.36 0.03 0.38 0.46 0.04
(e) StackedDiffusion [41] 0.43 0.02 0.42 0.57 0.07
(f) ShowHowTo 0.52 0.34 0.42 0.72 0.12
(g) Random 0.19 0.00 0.01 0.26 0.00
(h) Stable Diffusion [48$\dagger$ 0.70 0.03 0.44 0.84 0.03
(i) Copy of the input image 0.19 0.62 0.39 0.26 0.26
(j) Source sequences 0.50 1.00 0.56 0.60 1.00
@@
Table 1: Comparison with state-of-the-art on the ShowHowTo and the WikiHow datasets. Out of all the visual instruction generation methods, our method best follows the input prompts while being consistent with the input image.
Compared methods. We compare ShowHowTo to state-of-the-art methods for visual instruction generation as well as various baselines. For image-to-image methods (a-c), we generate the visual instructions sequence by iteratively using the last generated image as the input for the next step generation to achieve temporal consistency. (a) InstructPix2Pix [12] is trained to manipulate input images according to a text prompt by training on synthetic paired image data. In contrast, (b) AURORA [35] is trained on a manually curated dataset of image pairs from videos, while (c) GenHowTo [53] extracts the image pairs for training from instructional videos automatically.
Methods that generate image sequences (d-e) do not accept an input image, therefore, we apply the common input masking approach, where the first denoised frame of the sequence is replaced by the noised ground truth frame in each step of the generation. The (d) Phung et al. [45] method generates consistent sequences of visual instructions by attending to all frames in the sequence in the spatial attention layers. As it is a training-free method, we reimplement it and use it with the Stable Diffusion backbone [48]. (e) StackedDiffusion [41] is trained on WikiHow illustrated image sequences. It generates the image sequence as a single tiled image. Similarly to the related work, we evaluate all methods in zero-shot setup without finetuning.
@@
Quantitative results. We show the results of different methods on our ShowHowTo test set as well as on a subset of WikiHow sequences [62] in Table 1. The methods trained to perform localized edits (a-b) generate outputs fairly consistent with the input image (see the Scene Consistency metric), yet they fail to properly capture the instructions described by the text prompts (evidenced by the Step Faithfulness metric). On the other hand, methods for generating sequences of visual instructions (d-e) model the instructions well according to the input text prompts, but they perform poorly in scene consistency. In contrast, our method (f) generates visual instructions that are consistent with the input scene and correctly capture the action specified by the prompt. Our method even generates images that are more faithful to the input textual instructions than the dataset sequences (j) (see the Step Faithfulness metric). There are two reasons for this: objects in real images can appear small or occluded, impacting CLIP matching, and sometimes steps do not appear visually in the video.
Additionally, in Appendix E, we test our method in a zero-shot setup on the GenHowTo benchmark [53] and report additional quantitative metrics on the ShowHowTo dataset.
Figure 4: User study results. Win rates of the ShowHowTo method against baselines from pairwise forced decision user evaluations, divided into step, scene, and task. Values larger than 50% indicate ShowHowTo is preferred over the other methods (right).
User study. We present a user study with 9 participants evaluating sequences from 100 randomly sampled tasks from our test set. Each participant compared 50 ShowHowTo sequences with baselines using three criteria: (1) Step Faithfulness (Which sequence better follows the steps?), (2) Scene Consistency (Which sequence is more likely to come from the same video?), and (3) Task Faithfulness (Which sequence accurately depicts the instructions for the task of [task]?). As shown in Figure 4, ShowHowTo outperforms all baselines. Notably, users preferred our generations over sequences from source videos in 42% of cases for both step and scene metrics, which may be attributed to instructional videos not showing good views of steps at times and the high quality of our generation. Lower task faithfulness scores against the source sequences suggest room for improvement in future methods. More details are in Appendix D.
@@ 5.3 Ablations
We evaluate the key design decisions of our proposed method, i.e., the model conditioning, training data, and variable sequence length training, in the next paragraphs. Furthermore, additional performance analysis of the trained model is available Appendix E.
Text conditioning type Step Scene Task Average
@@
Text model conditioning. We evaluate how different types of text conditioning affect model performance. For video models, it is common to provide a single text prompt for conditioning. However, visual instructions vary substantially from one another, possibly requiring different conditioning. We construct a single prompt for each sequence by concatenating all step prompts and by summarizing the step prompts using an LLM [19]. In Table 2, we show that our choice of separate prompt per step significantly outperforms both of the single prompt variants. Additionally, we demonstrate that using the free-form step description for $\tau_{0}$ outperforms the fixed prompt ‘an image’.
Model training data Step Scene Task Average
@@
Training data. In Table 3, we analyze the impact of different training datasets by comparing our dataset with two related instructional datasets of similar scale and task coverage: (i) HowToStep [37], which contains automatically extracted video-text sequences from cooking videos, and (ii) WikiHow-VGSI [62], which consists of manually created image-text sequences from WikiHow articles, where the images primarily consist of digitally drawn illustrations. To train on the HowToStep dataset, we select the middle frame of each video segment as the visual instruction frame. We observe significantly worse performance caused both by the lack of precise instruction frame information as well as very noisy video segments (e.g., the dataset contains ‘Thank you for watching!’ segments which are not instructional). The performance is also worse when compared to the model trained only on the food-related ShowHowTo sequences that are extracted from the very same videos as the HowToStep sequences, indicating a superiority of our sequence extraction process. Training on WikiHow-VGSI yields higher Step Faithfulness score, likely due to the quality of manually matched images and prompts. However, the overall performance remains significantly below our approach, as the model primarily learns from illustrated images, resulting in less consistent scene generation.
Training sequence length Step Scene Task Average
@@
Variable sequence length. Instructions are often of variable length, therefore, one of the key model requirements is to support generating image sequences of different lengths. While attention-based architectures allow for any sequence length, the question is how to train such a model. We test training the model on variable sequences of up to 4 frames, 8 frames, and 16 frames. We also train the model on sequences of length 8 only. In Table 4, we show that training on shorter sequences up to 4 frames results in high Scene Consistency but low Step Correctness, while training on sequences up to 16 frames is the opposite. This can be attributed to the fact that short sequences often keep the same background across the whole sequence, encouraging the model to preserve the background at the expense of the prompt. Longer sequences, on the other hand, have more variation of the background, e.g., as the task moves from the counter to the hob. The model is thus less likely to enforce the background during inference. Lastly, we also show that it is important to train always on consecutive sequences of visual instructions. If a subset of visual instructions from a video is sampled randomly (with the temporal ordering preserved), the scene consistency is decreased (Table 4, row 3).
@@ 5.4 Qualitative Results
Qualitative results in Figure 1, Figure 6, and Figures 13, 14, and 15 in the appendix demonstrate the key strengths of the ShowHowTo model. It consistently preserves the scene as well as various objects, tools, and ingredients from the user-provided input image (e.g., pot and vegetables in Figure 6, first row). It dynamically adjusts the viewpoint to emphasize key actions. Similarly to vanilla text-to-image and text-to-video models, our model can also introduce plausible task-relevant objects (e.g., knives or bowls) if these objects are not present in the user-provided input image. Notably, the model effectively adapts human poses to demonstrate various object manipulations (e.g., flower arranging in Figure 6, third row).
Figure 5 shows qualitative comparisons with related methods. Methods for generating instructional sequences (StackedDiffusion [41] and Phung et al. [45]) fail to preserve the scene from the input image. For example, they generate blue, black, or wooden kitchen countertop instead of the glossy white one from the input image. On the other hand, image-to-image approaches (GenHowTo [53], AURORA [35], and InstructPix2Pix [12]) perform minimal edits and propagate errors through the output image sequence due to iterative generation (e.g., the persistent floating dough generated by the GenHowTo method).
Limitations. While our method can generate complex step-by-step visual instructions conditioned on the input image, it inherits the limitations of the models it is based on and introduces new limitations stemming from the novel source of training data. ShowHowTo model can struggle to maintain object states across many frames; for example, it can generate an image with raw meat after it was cooked in previous steps. Though the model often correctly generates common objects from instructional videos, for rare objects, such as electrical components, the model may generate objects in physically impossible configurations. Please see Figure 16 in the appendix for failure case examples.
@@ 6 Conclusion
This work explores, for the first time, generating environment-specific visual instructions to accomplish a user-defined task. We introduce a fully automated and scalable pipeline to create a dataset of 578K instructional image-text sequences from online videos, without requiring any manual supervision. Using this data, we train the ShowHowTo model to generate contextualized step-by-step visual instructions. Experiments demonstrate superior ability to generate accurate, scene-consistent instructional steps across various HowTo tasks, outperforming existing methods. We believe this work opens new avenues for personalized guidance in assistive technologies and step-by-step goal generation for robot planning.
Acknowledgements
We acknowledge VSB – Technical University of Ostrava, IT4Innovations National Supercomputing Center, Czech Republic, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (grant ID: 90254).
Research at the University of Bristol is supported by EPSRC UMPIRE (EP/T004991/1) and EPSRC PG Visual AI (EP/T028572/1). Prajwal Gatti is partially funded by an uncharitable donation from Adobe Research to the University of Bristol.
This research was co-funded by the European Union (ERC FRONTIER, No. 101097822 and ELIAS No. 101120237) and received the support of the EXA4MIND project, funded by the European Union’s Horizon Europe Research and Innovation Programme, under Grant Agreement N° 101092944. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.